# Ant Colony Sampling with GFlowNets for Combinatorial Optimization

### v1

@@ 4.1 MDP formulation
This subsection formulates a Markov decision process (MDP) by using TSP as an example, similar to the preliminaries. A TSP solution is obtained through the sequential decision-making process of GFlowNets as follows:
State $s$:
The initial state $s_{0}$ is an empty set. The state $s_{t}$ at $t>0$ is a sequence of previous actions: $(a_{0},\ldots,a_{t-1})$.
Action $a$:
The action $a_{t}$ is a selection of a node to visit at time $t$ among un-visited nodes: $a_{t}\in\{1,\ldots,N\}\setminus\{a_{0},\ldots,a_{t-1}\}$. If all nodes are visited, a deterministic terminating action is applied, which maps a final state $s_{N}$ into a solution, $x(\mathcal{G})$.
Reward $R(x(\mathcal{G});\beta)$:
The reward is defined as the unnormalized energy distribution of the solution $x(\mathcal{G})$, i.e.,
$R(x(\mathcal{G});\beta)=e^{-\beta\mathcal{E}(x(\mathcal{G}))}=e^{-\beta f(x(\mathcal{G}))}.$ (6)
Here, $\mathcal{E}(x(\mathcal{G}))$ is energy for solution $x(\mathcal{G})$, which is the same as the tour length of the solution for TSP, i.e., $\mathcal{E}(x(\mathcal{G}))=f(x(\mathcal{G}))$. The $\beta$ is the inverse energy temperature, which is a hyperparameter.
@@ 4.2 Policy Configuration
We parameterize the forward policy $P_{F}$ using a graph neural network (GNN) with parameter $\theta$ that maps an input graph instance $\mathcal{G}$ to the output edge feature $\eta(\mathcal{G};\theta)$, following Ye et al. (2023). Then, the forward policy when $t<N$ is expressed as follows:
$\begin{split}&P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G};\theta),\rho)\\ &\propto\begin{cases}\rho_{a_{t-1},a_{t}}\eta_{a_{t-1},a_{t}}(\mathcal{G};\theta)&\text{if $a_{t}\notin s_{t}$ }\\ 0&\text{otherwise}\end{cases}\end{split}$ (7)
Note that $\rho$ and $\eta(\mathcal{G};\theta)$ denote the pheromone trails and learned heuristic matrix in ACO.
Subsequently, we can define a trajectory from the initial state to a terminal solution $\tau=(s_{0},\ldots,s_{N},x(\mathcal{G}))$ and the forward probability of that trajectory:
$P_{F}(\tau;\eta(\mathcal{G};\theta),\rho)=\prod_{t=0}^{N-1}P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G};\theta),\rho)$ (8)
@@
For the backward policy, we set $P_{B}=\frac{1}{2N}$, emulating the maximum entropy GFlowNets (Zhang et al., 2022a) to make uniform exploration over symmetric trajectories which terminate into identical solution. Here, $P_{B}$ represents a uniform distribution across all $2N$ feasible trajectories leading to a solution $x(\mathcal{G})$. This is particularly pertinent in the context of the TSP involving $N$ cities, where $2N$ permutations culminate in the same Hamiltonian cycle $x(\mathcal{G})$. The count of symmetric trajectories depends on the task’s characteristics. For an exhaustive discussion on solution symmetry across various COPs, please refer to Appendix E.
@@ 4.3 Solution Sampling with GFACS
In the sampling process, we assume the existence of the learned GNN parameterized with $\theta^{*}$; see Section 5 for details of the training procedure. The learned GNN provides the heuristic matrix $\eta(\mathcal{G};\theta^{*})$ for ACO algorithm. Thus, the remaining process is to update pheromone parameter $\rho$ with solutions sampled by $P_{F}(\tau;\eta(\mathcal{G};\theta^{*}),\rho)$ during the solution searching process. As the $\eta(\mathcal{G};\theta^{*})$ is trained with GFlowNets to sample diverse high-rewarded (i.e., low energy) samples, $P_{F}$ generates good candidates for updating $\rho$. The overall process is illustrated in Figure 3.
5 Training GFACS
This section introduces how to train GFACS, especially the GNN component $\theta$, in the GFlowNet framework. GFACS is trained via iterations of the following two steps. First (Step A; Section 5.1), we collect experiences, each being a triplet of input instance graph $\mathcal{G}$, a complete trajectory $\tau_{\rightarrow x(\mathcal{G})}$, and the energy of the solution $\mathcal{E}(x(\mathcal{G}))$. We employ $\beta$ annealing and guided exploration for better exploration and energy reshaping for proper credit assignment. Next (Step B; Section 5.1), the parameterized forward policy $P_{F}$ is trained using the collected experiences. Here, we utilize a shared energy normalization to stabilize training.
5.1 Step A: Experience Collection
@@
For every iteration, we first sample $M$ problem instances from a random sampler $P(\mathcal{G})$, i.e., $\mathcal{G}_{1},\ldots,\mathcal{G}_{M}\sim P(\mathcal{G})$. Details of instance sampling are described in Appendix B.
Collecting exploration samples with $\beta$ annealing. For each problem instance $\mathcal{G}_{m}$, we collect $K$ exploration solutions by following Eq. 8, resulting in an exploration batch, $\mathcal{B}_{\text{explore}}=\{(\mathcal{G}_{m},\tau^{m,k}_{\rightarrow x_{m,k}},\mathcal{E}(x_{m,k})\}_{m=1,\ldots,M,k=1,\ldots,K}$. To encourage exploration for solution sampling, we introduce an annealing scheme for the inverse energy temperature $\beta$. Rather than set the $\beta$ to a fixed value, we gradually increase the beta toward the target value along the training procedure. This makes the policy in the early stage of training to be exploratory, preventing it from getting stuck into a local optimum.
Figure 4: Training of GFACS using GFlowNets and our novel training tricks of guided exploration and energy reshaping.
Collecting exploitation samples with guided exploration. To cope with the over-exploration problem of GFlowNet, we also collect $K$ exploitation samples for each $\mathcal{G}_{m}$ using guided exploration. We use a heuristic local search operator ($LS$), such as 2-opt for TSP, as external guidance. The local search operator directly refines a given solution into a locally optimized one, i.e., $x^{\prime}\leftarrow LS(x)$. We then use $P_{B}$ to collect a trajectory $\tau_{\rightarrow x^{\prime}}$ for each $x^{\prime}$, obtaining a batch of exploitation experiences, $\mathcal{B}_{\text{exploit}}=\{(\mathcal{G}_{m},\tau^{m,k}_{\rightarrow x^{\prime}_{m,k}},\mathcal{E}(x^{\prime}_{m,k})\}_{m=1,\ldots,M,k=1,\ldots,K}$.
However, a powerful local search operator may not exist for some problems, especially for less-explored ones. In such cases, we use the Top-$L$ guidance technique, described as follows. For a given problem instance $\mathcal{G}_{m}$, we first update the pheromone trails using the exploration samples collected before. After that, $L\times K$ new sample solutions are generated with the updated pheromone, where $L$ is a hyperparameter. Among the new samples, we select top $L$ solutions with the lowest energy and also select $(K-L)$ random solutions, resulting in $K$ exploitation samples for each $\mathcal{G}_{m}$, and thus obtaining $B_{\text{exploit}}$ like before.
@@
Energy reshaping for exploratory samples. In the case where we have a local search operator, we reshape the energy for the exploration $x$ in $\mathcal{B}_{\text{explore}}$ using the energy of the refined solution $x^{\prime}=LS(x)$ as follows:
$\mathcal{E}_{\text{reshaped}}(x)=\alpha\mathcal{E}(x^{\prime})+(1-\alpha)\mathcal{E}(x)$ (9)
This is to evaluate an underrated sample that initially had high energy (i.e., low reward) but can be easily changed to have low energy (i.e., high reward) after the local search (see Figure 2). This reshaping changes the reward values in $\mathcal{B}_{\text{explore}}$ accordingly, while $\mathcal{B}_{\text{exploit}}$ remains the same. Note that this reshaping cannot be applied for TopK guidance because $x^{\prime}$ cannot be determined solely with $x$ in that case.
@@
5.2 Step B: Training with Collected Experiences
@@
where ($\begin{split}&\mbox{where }\mathcal{L}(\theta;\mathcal{B})\\ &=\sum_{m=1}^{M}\sum_{k=1}^{K}\left(\log\frac{2NP_{F}(\tau^{m,k}_{\rightarrow x_{m,k}};\eta(\mathcal{G}_{m};\theta),\rho)Z_{\theta}(\mathcal{G}_{m})}{R(x_{m,k};\beta)}\right)^{2}.\end{split}$ (11)
The $2N$ comes from the fact that TSP has $2N$ permutations, as discussed in Section 4.2. In addition, our target sampling distribution is proportional to reward $p(x;\mathcal{G})\propto R(x(\mathcal{G}))$. However, in most COPs, including TSP, the scale of $R(x(\mathcal{G}))$ is significantly different, possibly leading to unstable training due to the high variance of rewards. Similarly to previous works that stabilize REINFORCE training (Kwon et al., 2020; Kim et al., 2022b), we normalize the energy $\mathcal{E}=-\log R$ using the sample mean over a $K$ samples for the same $\mathcal{G}_{m}$, i.e.,
@@
We call this technique a shared energy normalization. Note that the shared energy normalization is applied independently to the two batches, $\mathcal{B}_{\text{explore}}$ and $\mathcal{B}_{\text{exploit}}$, as the energy distribution between the two batches is often significantly different to each other, especially when a heuristic local search is applied.
Figure 5: Results of ACO algorithms on various CO tasks. Our GFACS outperforms every ACO baseline. The results are averaged over 3 independent models evaluated on the held-out test sets, and the shade indicates the min-max range of the 3 models. GFACS, without guided exploration, gives poorer performances. This verifies the effectiveness of our algorithm.
@@ 6 Experiments
This section presents the experimental findings of the GFACS framework. First, we test GFACS against established ACO baselines in various CO tasks. Subsequently, we explore the trade-offs between diversity and optimality in relation to the inverse energy temperature $\beta$. Next, an ablation study is conducted to evaluate the effectiveness of our proposed training techniques. Lastly, we compare GFACS with the RL-based combinatorial optimization methods in two representative routing problems, the TSP and the capacitated vehicle routing problem (CVRP).
6.1 Setup
Implementation and hyperparameters. GFACS is built upon the implementation established by DeepACO (Ye et al., 2023). As our methodology mainly focuses on improving training algorithms rather than network architecture, we simply adopt the GNN architecture proposed in DeepACO, except the additional parameters for partition function $Z_{\theta}$ in Eq. 11. For detailed network architecture and hyperparameters, please refer to Appendix A. One important hyperparameter that GFACS newly introduces is the inverse energy temperature $\beta$, described in Eq. 6. We offer an analysis of the impact and significance of $\beta$ in Section 6.4.
Test settings. Each algorithm is tested on the held-out datasets generated or obtained in advance. Unless otherwise stated, the results presented throughout this section are the averaged values from three independent models, each trained with a distinct random seed.
@@ 6.2 Comparison with ACO algorithms
Here, we compare GFACS against the classical ACO and the recent DeepACO algorithm across various CO tasks.
@@
Problems. Following the approach of Ye et al. (2023), we evaluate our algorithm in a range of COPs. This includes four routing problems – TSP, Capacitated Vehicle Routing Problem (CVRP), Prize-Collecting Vehicle Routing Problem (PCTSP), and Orienteering Problem (OP) – along with two scheduling problems – Sequence Ordering Problem (SOP) and Single Machine Total Weighted Tardiness Problem (SMTWTP) – and a grouping problem, the Bin Packing Problem (BPP). For detailed problem configurations, we direct readers to Appendix B or Ye et al. (2023) as we mostly follow their problem settings. For TSP and CVRP, we use local search algorithms with neural-guided perturbation following Ye et al. (2023). For other problems, we use TopK guidance as described in Section 5.1.
Baselines. We use ACO and DeepACO as baseline algorithms. Both baselines and GFACS use Ant System (Dorigo et al., 1996) for pheromone update (Eq. 2). We also test GFACS without guided exploration to analyze its effect on the performance. We use 100 ants and $T=100$ ACO iterations for all the algorithms.
Results. As shown in Figure 5, GFACS consistently surpasses both the classic ACO algorithm and DeepACO across all benchmark problems. Especially for routing problems, GFACS outperforms DeepACO by a considerable margin. This is particularly remarkable given that both algorithms utilize identical neural networks and local search operators, highlighting the crucial role of our training strategy. Without guided exploration, GFACS generally performs badly, likely due to the over-exploration problem of GFlowNets. This issue is effectively mitigated with the proposed guided exploration, showing the efficacy of our method. Interestingly, in CVRP, OP, and PCTSP, GFACS without guided exploration outperforms DeepACO after a certain number of ACO iterations. We believe that this is attributed to the increased solution diversity of GFACS, which comes from the generative nature of the policy and implied DAG structure that incorporates the solution symmetricity.
Figure 6: The trade-off between diversity and optimality as the inverse energy temperature $\beta$ changes. By adjusting $\beta$, we can control the trade-off of having a Pareto frontier to compare with DeepACO. The reported values are obtained with $T=1$, i.e., the solutions are not affected by the pheromone.
6.3 Analysis of $\beta$ and Diversity-Optimality Trade-off
In this subsection, we explored the diversity-optimality trade-off by varying the inverse energy temperature $\beta$ (Eq. 6) from low to high values. For analysis, we generate $K$ solution samples from policies trained with GFACS with different values of $\beta$ and DeepACO (with $\rho=1$ so that the pheromone does not affect the result). We use average pairwise Jaccard distance as a diversity metric (formally defined in Appendix D). Results for TSP200 (Figure 6) show that lower $\beta$ values lead to more diverse but less optimal solutions, while higher $\beta$ values result in less diversity but higher optimality. GFACS confirms the anticipated trade-off between diversity and optimality, establishing a Pareto frontier over DeepACO by offering more diverse and high-quality solutions. This advantage is particularly beneficial when an effective local search mechanism is available, increasing the likelihood of achieving superior local optima, as illustrated in Figure 2. However, excessively high $\beta$ (e.g., $2000$ in our result) degrades the optimality, supposedly because too high $\beta$ value prevents exploration during training, which in turn results in sub-optimality. This empirical evidence supports our design choice of $\beta$ annealing, which encourages exploration in the early stage of training.
Table 1: An ablation study examining various design elements of our method for TSP200 and CVRP200. Here, the optimality gap is used as a metric, where lower values indicate better performance.
Method TSP200 CVRP200
Ours 0.26% 1.09%
-- $\beta$ annealing 0.27% 1.11%
-- guided exloration 0.88% 1.42%
-- energy reshaping 0.59% 1.09%
-- shared energy normalization 3.11% 2.88%

### v2

@@ 4.1 MDP formulation
The solution construction procedure by GFlowNets can be modeled as a Markov decision process (MDP). We illustrate this procedure using the Traveling Salesman Problem (TSP) as an example. This formulation can be easily adapted to general combinatorial optimization problem with minor modifications.
State $s$: The initial state $s_{0}$ is an empty set. The state $s_{t}$ at $0<t<n$ is a partial combinatorial solution of prensented with previous actions: $\{a_{0},\ldots,a_{t-1}\}$ and $s_{n}$ is completed solution.
Action $a$: The action $a_{t}$ is a selection of a node to visit at time $t$ among un-visited nodes: $a_{t}\in\{1,\ldots,N\}\setminus\{a_{0},\ldots,a_{t-1}\}$. If all nodes are visited, a deterministic terminating action is applied, which maps a final state $s_{N}$ into a solution, $x(\mathcal{G})$.
Reward $R(x(\mathcal{G});\beta)$: The reward is defined as the unnormalized Boltzmann distribution of the solution $x(\mathcal{G})$ with energy $\mathcal{E}(x(\mathcal{G}))$. Specifically, the reward is given by $R(x(\mathcal{G});\beta)=e^{-\beta\mathcal{E}(x(\mathcal{G}))}$. Here, $\mathcal{E}(x(\mathcal{G}))$ represents the energy of the solution $x(\mathcal{G})$, which corresponds to metrics such as the tour length in the TSP.
The coefficient $\beta$, known as the inverse temperature, is a hyperparameter.
@@ 4.2 Policy configuration
We parameterize the forward policy $P_{F}$ using a graph neural network (GNN) with parameter $\theta$ that maps an input graph instance $\mathcal{G}$ to the output edge matrix $\eta(\mathcal{G};\theta)$, following Ye et al. [79]. Then, the forward policy when $t<N$ is expressed as follows:
$P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G};\theta),\rho)\propto\begin{cases}\rho_{a_{t-1},a_{t}}\eta_{a_{t-1},a_{t}}(\mathcal{G};\theta)&\text{if $a_{t}\notin s_{t}$ }\\ 0&\text{otherwise}\end{cases}$ (5)
Note that $\rho$ and $\eta(\mathcal{G};\theta)$ denote the pheromone trails and learned heuristic matrix in ACO. Subsequently, we can define a trajectory from the initial state to a terminal solution, with $\tau=(s_{0},\ldots,s_{N},x(\mathcal{G}))$, and the forward probability of that trajectory,
$P_{F}(\tau;\eta(\mathcal{G};\theta),\rho)=\prod_{t=0}^{N-1}P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G};\theta),\rho).$ (6)
@@
For the backward policy, we set $P_{B}=\frac{1}{2N}$, emulating the maximum entropy GFlowNets [83] to make uniform exploration over symmetric trajectories which terminate into identical solution. Here, $P_{B}$ represents a uniform distribution across all $2N$ feasible trajectories leading to a solution $x(\mathcal{G})$. This is particularly pertinent in the context of the TSP involving $N$ cities, where $2N$ permutations culminate in the same Hamiltonian cycle $x(\mathcal{G})$. The count of symmetric trajectories depends on the task’s characteristics. For an exhaustive discussion on solution symmetry across various COPs, please refer to Appendix E.
@@ 4.3 Solution sampling with GFACS
In the sampling process, we assume the existence of the learned GNN parameterized with $\theta^{*}$; see Section 5 for details of the training procedure.
@@
The learned GNN provides the heuristic matrix $\eta(\mathcal{G};\theta^{*})$ for ACO algorithm. Thus, the remaining process is to update pheromone parameter $\rho$ with solutions sampled by $P_{F}(\tau;\eta(\mathcal{G};\theta^{*}),\rho)$ during the solution searching process. As the $\eta(\mathcal{G};\theta^{*})$ is trained with GFlowNets to sample diverse high-rewarded (i.e., low energy) samples, $P_{F}$ generates good candidates for updating $\rho$. The overall process is illustrated in Figure 1.
5 Off-policy training of GFACS
This section introduces how to train GFACS, especially the GNN component parameterized with $\theta$, in the GFlowNet framework. GFACS is trained via iterations of the off-policy experience collection (Step A) and training (Step B).
5.1 Step A: Off-policy experience collection with local search and energy reshaping
Figure 2: Hindsight observation with local search and energy reshaping to control the trade-off between exploration and exploitation.
@@ 5.1 Step A: Off-policy experience collection with local search and energy reshaping
In this step, we collect experiences on given graph instances. The experiences in step A will be used for training GFACS in step B. For every iteration, we first sample $M$ problem instances from a random sampler $P(\mathcal{G})$, i.e., $\mathcal{G}_{1},\ldots,\mathcal{G}_{M}\sim P(\mathcal{G})$. Details of instance sampling are described in Appendix B. For each problem instance $\mathcal{G}_{m}$, we collect $K$ solutions by following Eq. 6 as $\mathcal{B}_{\text{explore}}=\{(\mathcal{G}_{m},\tau^{m,k}_{\rightarrow x_{m,k}},\mathcal{E}(x_{m,k})\}_{m=1,\ldots,M,k=1,\ldots,K}$.
To exploit high-reward regions in the large-scale combinatorial space, we also collect $K$ exploitation samples for each $\mathcal{G}_{m}$ using a local search operator ($LS$), such as 2-opt for TSP, as a powerful behavioral policy. The local search operator directly refines a given solution into a locally optimized one, i.e., $x^{\prime}\leftarrow LS(x)$. We then use the backward policy $P_{B}$ to sample a trajectory $\tau_{\rightarrow x^{\prime}}$ for each $x^{\prime}$, obtaining a batch of “exploitation experiences,” $\mathcal{B}_{\text{exploit}}=\{(\mathcal{G}_{m},\tau^{m,k}_{\rightarrow x^{\prime}_{m,k}},\mathcal{E}(x^{\prime}_{m,k})\}_{m=1,\ldots,M,k=1,\ldots,K}$. Please note that we also suggest a general-purpose local search operator that partially destroys a solution and then repairs for problems that lack domain-specific local search methods. Please refer to Appendix C for details about the destroy-and-repair local saerch.
While the concept of using local search for exploitation has been explored in recent GFlowNet works [39], our distinct approach involves utilizing the local search (LS) operation to facilitate hindsight observation—a forward-looking insight derived from extending the search beyond the current state. We focus on evaluating samples that were initially overlooked due to their high energy (i.e., low reward) but can be transformed to exhibit low energy (i.e., high reward) following the local search.
@@
To achieve this, we modify the energy (in a way we call energy reshaping) for the exploration around $x$ in $\mathcal{B}_{\text{explore}}$ by using the energy of the refined solution $x^{\prime}=LS(x)$ as follows:
$\mathcal{E}_{\text{reshaped}}(x)=\alpha\mathcal{E}(x^{\prime})+(1-\alpha)\mathcal{E}(x).$ (7)
@@
5.2 Step B: Training with collected experiences with shared energy normalization
@@
where ($\mbox{where }\mathcal{L}(\theta;\mathcal{B})=\sum_{m=1}^{M}\sum_{k=1}^{K}\left(\log\frac{2NP_{F}(\tau^{m,k}_{\rightarrow x_{m,k}};\eta(\mathcal{G}_{m};\theta),\rho)Z_{\theta}(\mathcal{G}_{m})}{R(x_{m,k};\beta)}\right)^{2}.$ (9)
The $2N$ comes from the fact that TSP has $2N$ permutations, as discussed in Section 4.2. In addition, our target sampling distribution is proportional to reward $p(x;\mathcal{G})\propto R(x(\mathcal{G}))$. However, in most COPs, including TSP, the scale of $R(x(\mathcal{G}))$ can vary significantly across instances, possibly leading to unstable training due to the high variance of logarithms of rewards. Similarly to previous works that stabilize REINFORCE training [43, 37], we normalize the energy $\mathcal{E}=-\log R$ using the sample mean over $K$ samples for the same $\mathcal{G}_{m}$, i.e.,
@@
We call this technique a shared energy normalization. Note that the shared energy normalization is applied independently to the two batches, $\mathcal{B}_{\text{explore}}$ and $\mathcal{B}_{\text{exploit}}$, as the energy distribution between the two batches is often significantly different from each other, especially when a heuristic local search is applied. To this end, the gradient loss is computed as follows:
$\nabla_{\theta}\mathcal{L(\theta)}\approx\nabla_{\theta}\sum_{m=1}^{2M}\sum_{i=1}^{I}\left(\log\frac{2NP_{F}(\tau^{i,m}_{\rightarrow x_{i,m}};\eta(\mathcal{G}_{m};\theta),\rho)Z_{\theta}(\mathcal{G}_{m};\theta)}{\tilde{R}(x_{i,m})}\right)^{2}.$ (11)
@@ 6 Experiments
This section presents experimental results to validate our algorithm [^1] Source code is available: https://github.com/ai4co/gfacs. . We compare GFACS with other ACO algorithms, and with competitive reinforcement learning algorithms for vehicle routing problems. Moreover, we conduct an ablation study, and perform an empirical comparison with previous GFlowNets training methods, which serve a similar purpose to our method. Below, we provide a brief overview of the experimental setup.
Implementation and hyperparameters. GFACS is built upon the implementation established by DeepACO [79]. As our methodology mainly focuses on improving training algorithms rather than network architecture, we simply adopt the GNN architecture proposed in DeepACO, except the additional parameters for partition function $Z_{\theta}$ in Equation 9. For detailed network architecture, hyperparameters and hardware settings, please refer to Appendix A. It is worth noting that one important hyperparameter that GFACS newly introduces is the inverse energy temperature $\beta$, which enables us to control the diversity-optimality trade-offs (see Section D.2 for detailed analysis). In addition, we employ Ant System [14] for the pheromone updates throughout the experiments, while we also provide experimental results with other ACO variants in Section D.3.
Test settings. Each algorithm is tested on the held-out datasets, each with 100 instances generated in advance except TSP, where we obtain the TSP test dataset with 128 instances from Ye et al. [79]. Unless otherwise stated, the results presented throughout this section are the averaged values from three independent models, each trained with a distinct random seed.
@@ 6.1 Comparison with ACO algorithms
Here, we compare GFACS against the vanilla ACO with and the DeepACO [79] algorithm across various CO tasks.
@@
Problems. Following the approach of Ye et al. [79], we evaluate our algorithm in a range of COPs. This includes four routing problems – TSP, Capacitated Vehicle Routing Problem (CVRP), CVRP with Time Window (CVRPTW), Prize-Collecting Vehicle Routing Problem (PCTSP), and Orienteering Problem (OP) – along with a scheduling problem – Single Machine Total Weighted Tardiness Problem (SMTWTP) – and a grouping problem, the Bin Packing Problem (BPP). For detailed problem configurations, we direct readers to Appendix B. For TSP, CVRP and CVRPTW, we use local search algorithms with perturbation similar to Ye et al. [79]. For others, we use the destroy-and-repair method (see Appendix C for more details).
Figure 3: Results of ACO algorithms on various CO tasks. Our GFACS outperforms every ACO baseline. The results are averaged over 3 independent models evaluated on the held-out test sets, and the shade indicates the min-max range of the 3 models.
Baselines. We use vanilla ACO with a human-designed heuristic matrix, DeepACO as baseline algorithms. We also test GFACS without the proposed off-policy training scheme (called On-policy TB). For inference, we use $K=100$ ants and $T=100$ ACO iterations for all the algorithms.
Results. Figure 3 show GFACS consistently outperforms both the classic ACO algorithm and DeepACO across all benchmark problems. Notably, for routing problems, GFACS surpasses DeepACO by a significant margin. This is particularly remarkable given that both algorithms use identical neural networks and local search operators, underscoring the importance of our training strategy. Additionally, the proposed off-policy training method enhances GFACS performance, demonstrating its effectiveness for GFlowNets.
In this section, we present experimental results comparing our method with reinforcement learning (RL)-based approaches, as both aim to develop combinatorial solvers without the need for labeled optimal data points, unlike supervised learning methods (e.g., [16, 70]).
Baselines. For TSP baselines, we select the AM [41], POMO [43], Sym-NCO [37] and Pointerformer [28] as RL-based constructive methods, DIMES [62] as RL-based heatmap-based methods, the Select and Optimize (SO) [9] as RL-based improvement methods. We also applied two search algorithms, the Efficient Active Search (EAS) [23] and Simulation Guided Beam Search (SGBS) [10], to POMO for a more thorough comparison. For CVRP, [^2] We also adopt TAM as a CVRP baseline; however, due to the inaccessibility of source code, we compare it using the reported result in the paper [24]. See Section D.5. we employ POMO and Sym-NCO as baselines with EAS and SGBS. Note that we restrict the duration of these searches to no more than 5 and 10 minutes for problems with 500 and 1000 nodes, respectively. We use Concorde [2] for TSP and PyVRP [77] for CVRP as oracle heuristic solvers to obtain near-optimal solutions that serve as baselines with $0.00\%$ gap. For ACO algorithms, we use 100 ants with $T=10$ with local search algorithms similar to Ye et al. [79].
Results. Among the baseline methods evaluated, our GFACS algorithm consistently delivers either superior or highly competitive performances. When compared to faster methods like AM or POMO, GFACS demonstrates substantially better results. Even after an additional search with EAS or SGBS, they fail to match GFACS’s performance. Compared to algorithms like DIMES and Pointformer, which operate at similar speeds, GFACS achieves improved performance. GFACS performs slightly worse than the oracle heuristic solver LKH-3, but offers an advantage of faster inference, especially in CVRP. Therefore, we can conclude that GFACS, as a general-purpose algorithm, holds significant appeal in the highly competitive benchmarks of both TSP and CVRP.
Table 1: Experimental results on TSP and CVRP. Results for methods with * are drawn from Ye et al. [79], Cheng et al. [8], and Jin et al. [28]. The reported objective value is averaged over 128 instances of TSP and 100 instances of CVRP. We report the average computation time per instance.
TSP500 TSP1000 CVRP500 CVRP1000
Method Obj. Gap(%) Time Obj. Gap(%) Time Method Obj. Gap(%) Time Obj. Gap(%) Time
Concorde 16.55 - 10.7s 23.12 - 108s PyVRP 62.96 - 21.0m 119.29 - 1.4h
LKH-3 16.55 0.02 31.7s 23.14 0.10 102s LKH-3 63.07 0.18 18.6m 119.52 0.20 2.1h
AM* 21.46 29.07 0.6s 33.55 45.10 1.5s
POMO* 20.57 24.40 0.6s 32.90 42.30 4.1s POMO 68.69 9.08 2s 145.74 22.14 3s
    +EAS 18.25 10.29 300s 30.77 33.06 600s     +EAS 64.76 2.85 5.0m 126.25 5.83 10.0m
    +SGBS 18.86 13.98 318s 28.58 23.61 605s     +SGBS 65.44 3.93 5.2m 127.90 7.21 10.1m
DIMES* 17.01 2.78 11s 24.45 5.75 32s SymNCO 68.81 9.28 2s 141.82 18.86 3s
SO* 16.94 2.40 15s 23.77 2.80 26s     +EAS 64.63 2.65 5.0m 125.58 5.27 10.0m
Pointerformer* 17.14 3.56 14s 24.80 7.90 40s     +SGBS 65.40 3.87 5.2m 127.53 6.90 10.1m
ACO 17.38 5.04 14s 24.76 7.09 57s ACO 64.62 2.63 36s 122.82 2.96 1.3m
DeepACO 16.84 1.77 15s 23.78 2.87 66s DeepACO 64.49 2.43 36s 122.15 2.40 1.3m
GFACS (ours) 16.80 1.56 15s 23.72 2.63 66s GFACS (ours) 64.19 1.95 36s 121.80 2.11 1.3m
Table 2: Experimental results on real-world datasets, TSPLib and CVRPLib-X. ‘#’ denotes the number of instances in each set. We report the average optimality gap using the best-known costs.
TSPLib # ACO DeepACO Ours CVRPLib-X # ACO DeepACO Ours
100-299 30 1.71% 1.25% 1.21% 100-299 43 2.50% 3.27% 2.69%
300-699 10 4.26% 2.70% 2.60% 300-699 40 3.75% 4.22% 3.71%
700-1499 12 7.01% 4.08% 3.88% 700-1001 17 4.64% 5.07% 4.32%
Evaluation on real-world datasets. We compare the performance of our model against baseline models on real-world TSP and CVRP instances from the TSPLib [64] and CVRPLib [72] X datasets. The models were trained on random uniform instances of TSP/CVRP with sizes 200 for dataset sizes 100-299, 500 for sizes 300-699, and 1000 for sizes 700-1499, with all other settings remaining the same. Table 2 presents the results. GFACS outperforms the baselines in TSPLib and large CVRPLib instances but slightly underperforms compared to vanilla ACO on small CVRPLib instances, where ACO with Swap* local search suffices for small-scale CVRP. As the problem scale increases, ACO’s performance declines, underscoring the importance of effective priors using GFlowNets for efficient pheromone posterior updates in large combinatorial spaces.